# Membrane phononic integrated circuits

### v1

@@ Abstract
Phononic circuits constructed from high tensile stress membranes offer a range of desirable features such as high acoustic confinement, controllable nonlinearities, low mass, compact footprint, and ease of fabrication. This tutorial presents a systematic approach to modelling and designing phononic integrated circuits on this platform, beginning with acoustic confinement, wave propagation and dispersion, mechanical and actuation nonlinearities, as well as resonator dynamics. By adapting coupled mode theory from optoelectronics to suspended membranes, and validating this theory with several numerical techniques (finite element modelling, finite difference time domain simulations, and the transfer matrix method), we then provide a comprehensive framework to engineer a broad variety of phononic circuit building blocks. As illustrative examples, we describe the implementation of several acoustic circuit elements including resonant and non-resonant variable-ratio power splitters, mode converters, mode (de)multiplexers, and in-line Fabry-Pérot cavities based on evanescent tunnel barriers. These building blocks lay the foundation for phononic integrated circuits with applications in sensing, acoustic signal processing, and power-efficient and radiation-hard computing.
@@ I The wave equation
While membranes under tensile stress can be built from a variety of amorphous materials, chiefly silicon nitride [25, 15, 26, 27] and silicon carbide [28, 29], and even strained crystalline materials such as graphene [30, 31] and silicon [32], this tutorial refers mainly to highly stressed thin-film silicon nitride membranes like those used in our previous work [12, 22, 33, 20]. Typical parameters for this material are $\rho=3200\,\mathrm{kg\,m^{-3}}$ and $\sigma\simeq 1\,\mathrm{GPa}$, giving $c\simeq 560\,\mathrm{m\,s^{-1}}$. Appendix A lists further relevant physical parameters.
@@ I.1 Rigidity vs tension justification
where $k$ is the wavenumber. From Eq. (3) we can determine that in thin film silicon nitride, the transition from membrane-like to plate-like behaviour occurs at frequencies in excess of hundreds of megahertz. As shown in Fig. 3(c), for an $80\,\mathrm{nm}$-thick membrane (as used in our previous work [12, 22, 20]) with tensile stress of $1\,\mathrm{GPa}$, the tension is an order of magnitude stronger than the rigidity until around $70\,\mathrm{MHz}$ and becomes comparable around $200\,\mathrm{MHz}$. For a thinner $20\,\mathrm{nm}$-thick membrane (as demonstrated in [37]) the corresponding values are $300\,\mathrm{MHz}$ and $900\,\mathrm{MHz}$ (which corresponds to a wavelength of $\sim 0.6\,\mathrm{\upmu m}$). This tutorial considers $10-100\,\mathrm{\upmu m}$ sized silicon nitride membranes which have typical operating frequencies $50\,\mathrm{MHz}$, so we will henceforth neglect the rigidity term in Eq. (2).
@@ II Membrane device nomenclature
Resonators, waveguides, and couplers can often be distinguished by their geometry—for example, waveguides tend to have larger aspect ratios than resonators, and couplers tend to have smaller lateral dimension than waveguides—but is better to categorise phononic components by their function, as the fabrication and design scope for phononic circuits is very large.
Figure 5: Pictoral example of a resonator (green), coupler (orange), and waveguide (blue). Circles illustrate the vibrational mode shapes, with the resonator and waveguide oscillating in their fundamental modes. The waveguide is comparatively much longer (indicated by the break).
@@ III Waveguides and dispersion
One of the key tasks in a phononic circuit is transporting and directing the flow of acoustic energy. To do this with we use phononic waveguides, which are the acoustic equivalent to wires in an electrical circuit or optical waveguides in a photonic circuit [43]. They are typically characterised by low transmission loss and large aspect ratios, and are used to route and store travelling waves—both classical and quantum [44].
Some of the material platforms used for phononic waveguides include index-contrast slabs [3, 45], phononic crystals [46, 47, 48, 49, 50, 51], and topological waveguides [52, 53, 54]. In comparison with these platforms, suspended membrane waveguides have highly desirable properties such as low loss rates and a highly manipulable dispersion relationship [11, 9, 12].
@@
This describes a wave of frequency $\Omega$ and wavenumber $k_{y}$ moving in the +y+y direction, with a characteristic lateral modeshape $\psi(x)$. $Z_{0}$ is the maximum out-of-plane ($z$ axis) amplitude of oscillation and contains the dimensions of length, such that $\psi(x)\in[-1,1]$ and is dimensionless.
@@
The left and right hands of the equation contain different variables and so must equal a constant. The only physically meaningful choice of constant is $k_{x}^{2}$, as then the solution for $\psi(x)$ is simply:
@@
The clamped edges boundary condition from Section I.2 require that $\psi(0)=\psi(W)=0$. This immediately implies $B=0$, and because $\psi(x)\in[-1,1]$ we also conclude $A=1$. Finally, the condition $\psi(W)=0$ implies that $k_{x}=n\pi/W$ for integral $n$. The dispersion relationship then becomes:
@@
Figure 7: (a) Dispersion relationship $\Omega(k)$ obtained from Eq. (9), plotted for the first four modes. (b) Example displacements for different frequencies, found by summing over the modes that are above cutoff frequency with equal amplitude and random phase in each mode. (c) Phase velocity $v_{p}$ obtained from Eq. (11). Dashed line indicates $c=\sqrt{\sigma/\rho}$. (d) Group velocity $v_{g}$ obtained from Eq. (12). Dashed line also indicates $c$. All values are calculated for a silicon nitride waveguide of width $W=30\,\mathrm{\upmu m}$, density $\rho=3200\,\mathrm{kg\,m^{-3}}$, and stress $\sigma=1\,\mathrm{GPa}$.
@@
Subfigures 7(c-d) illustrates how the phase velocity and group velocity depend on the wave frequency. There are two regimes worth noting. As can be seen from Equations (11) and (12), as $k\rightarrow 0$ and $\Omega\rightarrow\Omega_{c,n}$ from above, $v_{p}\rightarrow\infty$ and $v_{g}\rightarrow 0$. In this situation the length of the waveguide is similar to the wavelength and it oscillates like a very large aspect ratio resonator. Since $v_{g}$ describes the speed at which energy is transferred in the waveguide [58], we have recovered the fact that no energy is transported (across long distances) when the waveguide is excited below the cutoff frequency. Conversely, as $k\rightarrow\infty$, the phase velocity and group velocity both converge to $c=\sqrt{\sigma/\rho}.$ This is because the clamped-edges boundary conditions are less relevant in the context of short wavelengths. The situation instead resembles the propagation of plane waves in a spatially infinite membrane, with the dispersionless plane wave dispersion relation: $\Omega(k)=ck$.
@@ IV.1 Resonator eigenmodes
Whereas waveguides are bounded in one dimension, resonators have two dimensional boundary conditions. As we did in the case of waveguides, we can look for a separable solution for the resonator displacement that separates the time and spatial variables. Unlike the waveguide case, we can’t assume the solution will take the form of a travelling wave, so we will keep the $x$ and $y$ variables together:
@@
Equation (17) neglects dissipative and nonlinear processes, which introduce time dynamics beyond simple harmonic oscillation. It is possible to numerically solving two-dimensional continuum models with dissipation and nonlinearity, but this can be computationally expensive. However, a strength of Eq. (17) is that it suggests we can model some regions of a two-dimensional material as systems of one-dimensional oscillators. This is the principle behind the lumped-element model, which we introduce in the next section.
@@ IV.2 Lumped element model
where $z(t)$ describes the time dynamics and $\psi(x,y)$ describes the eigen modeshape. $z(t)$ has dimensions of length and $\psi(x,y)$ is dimensionless $\max_{x,y}|\psi(x,y)|=1$. Comparing this with Eq. (17), we have restricted attention to one eigenmode (eliminating $n$), merged the amplitude $Z_{0}$ with the time dynamics, and no longer assume the time dynamics correspond to harmonic oscillations.
@@
$\max(E_{\mathrm{kin}}^{\mathrm{wave}})$ $=\frac{1}{2}\rho h\iint\left(\frac{\partial}{\partial t}z(t)\psi(x,y)\right)^{2}\,\mathrm{d}x\mathrm{d}y$
@@
We can obtain the effective spring constant, $k_{\mathrm{eff}}=\Omega_{0}^{2}m_{\mathrm{eff}}$, from the elastic potential energy:
@@
In this derivation, Eq. (25) uses the two-dimensional version of the potential energy density in a string [24]. Equation (26) can be interpreted as the fact that the tensile restoring force acts to reduce curvature in the membrane. A mode with greater curvature therefore has a stronger effective spring constant.
@@
This equation constrains $\psi$ and provides some intuition for the relationship between modeshape and eigenfrequency. The numerator scales with increasing curvature and the denominator increases with mode area. This corresponds to the fact that eigenmodes confined to smaller areas with sharper curvature have higher eigenfrequencies.
@@ IV.3 Damping
$m_{\mathrm{eff}}\frac{\mathrm{d}^{2}z(t)}{\mathrm{d}t^{2}}+\Gamma_{\mathrm{eff}}\frac{\mathrm{d}z(t)}{\mathrm{d}t}+k_{\mathrm{eff}}z=F(t).$ (28)
Here, $\Gamma_{\mathrm{eff}}$ is the effective damping force coefficient, $F(t)$ is the drive force and $\Omega_{d}$ is the drive frequency. It is worth reiterating that all of the effective quantities are defined with respect to a single eigenmode.
@@ IV.4 Spectral response and quality factor
$m_{\mathrm{eff}}\frac{\mathrm{d}^{2}z(t)}{\mathrm{d}t^{2}}+m_{\mathrm{eff}}\gamma_{\mathrm{eff}}\frac{\mathrm{d}z(t)}{\mathrm{d}t}+k_{\mathrm{eff}}u(t)=F_{0}e^{i\Omega_{d}t}.$ (33)
@@
$\arg(Z_{0})=\arctan\left(\frac{\frac{\gamma}{\Omega_{0}}\frac{\Omega_{d}}{\Omega_{0}}}{1-\left(\frac{\Omega_{d}}{\Omega_{0}}\right)^{2}}\right).$ (35)
@@
The quality factor also equals the ratio of total energy stored in the oscillator, $E$, to the the energy lost per oscillation, $\Delta E$:
@@
The equivalence of Eq.s (37) and (38) can be shown by noting that the energy in the oscillator is proportional to maximum amplitude squared, and under small damping the maximum amplitude decays exponentially as described in Eq. (32).
@@
Figure 10: Experimental demonstration of how resonances correspond to distinct eigenmodes, using the example of a phononic error correcting device [62]. (a) Optical microscope image of a phononic error-correcting device which consists of three coupled silicon nitride resonators. The resonators are $100\,\mathrm{\upmu m}$ diameter circular drums, coupled with evanescent tunnels. The device in this image has tunnels of length $20\,\mathrm{\upmu m}$. (b) Spectral response from optical Doppler interferometry, measured on the bottom left resonator when driving on the top resonator. These data were taken on a device with $11\,\mathrm{\upmu m}$ tunnels. (c) Experimentally obtained images of the modeshapes labelled in subfigure (b), on the same device and again driving the top resonator. These images were taken by measuring the acoustic power spectral density around the drive frequency while driving at a particular resonance frequency. Higher-order drum modes can be observed by the shape of their node lines (locations where the amplitude of vibration is zero).
In this section we have considered just one resonance with effective parameters $k_{\mathrm{eff}}$, $m_{\mathrm{eff}}$, and $\gamma$. In practice a membrane will usually exhibit a range of resonances corresponding to the different eigenmodes. Fig. 10 shows an experimental example of this, where on a single device we can measure a large number of resonances, each corresponding to a particular spatial mode that can be experimentally observed [62]. In Fig. 10(b) the resonance peak heights and widths vary because the different modes are more or less efficiently actuated by the electrode, and have larger or smaller damping rates. Differences in actuation efficiency can be observed for the mode labelled ‘C’ in Fig. 10(c). Section VII goes into more detail about how the mode shape affects actuation efficiency and loss.
@@ V Duffing nonlinearity
Nonlinearity enables many of the more useful and complex behaviours in phononic devices—for example it is essential for solitons and frequency combs [63, 64, 65, 66, 67], logic [19, 18, 68, 69, 20], and other phenomena such as four-wave mixing [70, 36, 71, 18] and bifurcations [72, 63]. In suspended membranes two key sources of nonlinearity are elongation stresses during large out-of-plane vibration, and electrostatic nonlinearities. In both cases we can effectively model the nonlinearity by adding a quartic potential to the mechanical potential energy (we show this later). This model is called a Duffing nonlinearity, named after Georg Duffing (1861-1944) [73, 74].
@@
where $\alpha_{\mathrm{eff}}$ is the Duffing coefficient. Fig. 11 plots this model potential. When $\alpha_{\mathrm{eff}}$ is positive the nonlinearity is stiffening and increases the steepness of the potential walls. When $\alpha_{\mathrm{eff}}$ is negative the nonlinearity is softening nonlinearity and decreases the wall steepness.
@@
With Duffing nonlinearity the resonance frequency depends on the amplitude of oscillation. For ‘stiffening’ $\alpha_{\mathrm{eff}}>0$ the membrane will increase in resonance frequency at larger amplitudes. For ‘softening’ $\alpha_{\mathrm{eff}}<0$ the membrane will decrease in resonance frequency at larger amplitudes.
The shifting resonance frequency of a Duffing oscillator has widespread implications. It limits the dynamic range of a frequency-shift based nanomechanical sensor [75], and it complicates the problem of coupling multiple nonlinear oscillators. It also introduces new functionalities, such as amplitude bistability which we will look at shortly in Section V.2.
To incorporate Duffing nonlinearity into the lumped element model, we note from Eq. (39) that the restoring force $F_{\mathrm{res}}$ is altered from Hooke’s law to $F_{\mathrm{res}}=\mathrm{d}/U\mathrm{d}z=k_{\mathrm{eff}}z+\alpha_{\mathrm{eff}}z^{3}$. Substituting the additional term into the lumped element dynamical equation (Eq. (28)) with a sinusoidal drive, we obtain the Duffing equation:
@@
Equation (44) quantifies the response of the resonator to the applied driving force. Figure 12(a) illustrates this response. As predicted in Eq. (41) we see the resonance frequency shifts depending on the amplitude of vibration, with the direction of the shift corresponding to the sign of $\alpha_{\mathrm{eff}}$. The peak response increases for softening nonlinearities ($\alpha_{\mathrm{eff}}<0$), because to first order a softening nonlinearity reduces the effective spring constant of the oscillator (even as the force $F_{0}$ remains constant).

### v2

@@ Abstract
Phononic circuits constructed from high tensile stress membranes offer a range of desirable features such as high acoustic confinement, controllable nonlinearities, low mass, compact footprint, and ease of fabrication. This tutorial presents a systematic approach to modelling and designing phononic integrated circuits on this platform, beginning with acoustic confinement, wave propagation and dispersion, mechanical and actuation nonlinearities, as well as resonator dynamics. By adapting coupled mode theory from optoelectronics to suspended membranes, and validating this theory with several numerical techniques (finite element modelling, finite difference time domain simulations, and the transfer matrix method), we provide a comprehensive framework to engineer a broad variety of phononic circuit building blocks. As illustrative examples, we describe the implementation of several acoustic circuit elements including resonant and non-resonant variable-ratio power splitters, mode converters, mode (de)multiplexers, and in-line Fabry-Pérot cavities based on evanescent tunnel barriers. These building blocks lay the foundation for phononic integrated circuits with applications in sensing, acoustic signal processing, and power-efficient and radiation-hard computing.
@@ I The wave equation
While membranes under tensile stress can be built from a variety of amorphous materials, chiefly silicon nitride [25, 15, 26, 27] and silicon carbide [28, 29], and even strained crystalline materials such as graphene [30, 31] and silicon [32], this tutorial refers mainly to highly stressed thin-film silicon nitride membranes like those used in our previous work [12, 22, 33, 20]. Typical parameters for this material are $\rho=3200\,\mathrm{kg\cdot m^{-3}}$ and $\sigma\simeq 1\,\mathrm{GPa}$, giving $c\simeq 560\,\mathrm{m\cdot s^{-1}}$. Appendix A lists further relevant physical parameters.
@@ I.1 Rigidity vs tension justification
where $k$ is the wavenumber. From Eq. (3) we can determine that in thin film silicon nitride, the transition from membrane-like to plate-like behaviour occurs at frequencies in excess of hundreds of megahertz. As shown in Fig. 3(c), for an $80\,\mathrm{nm}$-thick membrane (as used in our previous work [12, 22, 20]) with tensile stress of $1\,\mathrm{GPa}$, the tension is an order of magnitude stronger than the rigidity until around $70\,\mathrm{MHz}$ and becomes comparable around $200\,\mathrm{MHz}$. A thinner $20\,\mathrm{nm}$-thick membrane (as demonstrated in [37]) the corresponding values are $300\,\mathrm{MHz}$ and $900\,\mathrm{MHz}$ (which corresponds to a wavelength of $\sim 0.6\,\mathrm{\upmu m}$). This tutorial consider $10-100\,\mathrm{\upmu m}$ sized silicon nitride membranes which have operating frequencies $100\,\mathrm{MHz}$, so we will henceforth neglect the rigidity term in Eq. (2).
@@ II Membrane device nomenclature
Resonators, waveguides, and couplers can often be distinguished by their geometry—for example, waveguides tend to have larger aspect ratios than resonators, and couplers tend to have smaller lateral dimension than waveguides—but it is better to categorise phononic components by their function, as the fabrication and design scope for phononic circuits is very large.
Figure 5: Pictorial example of a resonator (green), coupler (orange), and waveguide (blue). Circles illustrate the vibrational mode shapes, with the resonator and waveguide oscillating in their fundamental modes. The waveguide is comparatively much longer (indicated by the break).
@@ III Waveguides and dispersion
One of the key tasks in a phononic circuit is transporting and directing the flow of acoustic energy. To do this we use phononic waveguides, which are the acoustic equivalent to wires in an electrical circuit or optical waveguides in a photonic circuit [43]. They are typically characterised by low transmission loss and large aspect ratios, and are used to route and store travelling waves—both classical and quantum [44, 45, 46].
Some of the material platforms used for phononic waveguides include index-contrast slabs [3, 47, 5, 48, 49, 50], phononic crystals [51, 52, 53, 54, 55, 56], and topological waveguides [57, 58, 59]. In comparison with these platforms, suspended membrane waveguides have highly desirable properties such as low loss rates and a highly manipulable dispersion relationship [11, 9, 12, 60, 61].
@@
This describes a wave of frequency $\Omega$ and wavenumber $k_{y}$ moving in the +y+y direction, with a characteristic lateral modeshape $\psi(x)$. $Z_{0}$ is the maximum out-of-plane ($z$ axis) amplitude of oscillation and contains the dimensions of length, such that $\psi(x,y)$ is dimensionless and $\max(|\psi(x,y)|)=1$.
@@
The left- and right-hands of the equation contain different variables and so must equal a constant. The only physically meaningful choice of constant is $k_{x}^{2}$, as then the solution for $\psi(x)$ is simply:
@@
The clamped edges boundary condition from Section I.2 require that $\psi(0)=\psi(W)=0$. This immediately implies $B=0$, and because $\max(|\psi(x,y)|)=1$ we also conclude $A=1$. Finally, the condition $\psi(W)=0$ implies that $k_{x}=n\pi/W$ for integral $n$. The dispersion relationship then becomes:
@@
Figure 7: (a) Dispersion relationship $\Omega(k)$ obtained from Eq. (9), plotted for the first four modes. (b) Example displacements for different frequencies, found by summing over the modes that are above cutoff frequency with equal amplitude and random phase in each mode. (c) Phase velocity $v_{p}$ obtained from Eq. (11). Dashed line indicates $c=\sqrt{\sigma/\rho}$. (d) Group velocity $v_{g}$ obtained from Eq. (12). Dashed line also indicates $c$. All values are calculated for a silicon nitride waveguide of width $W=30\,\mathrm{\upmu m}$, density $\rho=3200\,\mathrm{kg\cdot m^{-3}}$, and stress $\sigma=1\,\mathrm{GPa}$.
@@
Subfigures 7(c-d) illustrate how the phase velocity and group velocity depend on the wave frequency. There are two regimes worth noting. As can be seen from Equations (11) and (12), as $k\rightarrow 0$ and $\Omega\rightarrow\Omega_{c,n}$ from above, $v_{p}\rightarrow\infty$ and $v_{g}\rightarrow 0$. In this situation the length of the waveguide is similar to the wavelength and it oscillates like a very large aspect ratio resonator. Since $v_{g}$ describes the speed at which energy is transferred in the waveguide [65], we have recovered the fact that no energy is transported (across long distances) when the waveguide is excited below the cutoff frequency. Conversely, as $k\rightarrow\infty$, the phase velocity and group velocity both converge to $c=\sqrt{\sigma/\rho}.$ This is because the clamped-edges boundary conditions are less relevant in the context of short wavelengths. The situation instead resembles the propagation of plane waves in a spatially infinite membrane, with the dispersionless plane wave dispersion relation: $\Omega(k)=ck$.
@@ IV.1 Resonator eigenmodes
Whereas waveguides are bounded in one dimension, resonators have two dimensional boundary conditions. As we did in the case of waveguides, we can look for a separable solution for the resonator displacement that separates the time and spatial variables. Unlike the waveguide case, we cannot assume the solution will take the form of a travelling wave, so we will keep the $x$ and $y$ variables together:
@@
Equation (17) neglects dissipative and nonlinear processes, which introduce time dynamics beyond simple harmonic oscillation. It is possible to numerically solve two-dimensional continuum models with dissipation and nonlinearity, but this can be computationally expensive. However, a strength of Eq. (17) is that it suggests we can model some regions of a two-dimensional material as systems of one-dimensional oscillators. This is the principle behind the lumped-element model, which we introduce in the next section.
@@ IV.2 Lumped element model
where $z(t)$ describes the time dynamics and $\psi(x,y)$ describes the eigenmode shape. $z(t)$ has dimensions of length and $\psi(x,y)$ is dimensionless $\max_{x,y}|\psi(x,y)|=1$. Comparing this with Eq. (17), we have restricted our attention to one eigenmode, merged the amplitude $Z_{0}$ with the time dynamics, and no longer assume the time dynamics correspond to harmonic oscillations.
@@
$\max(E_{\mathrm{kin}}^{\mathrm{wave}})$ $=\frac{1}{2}\rho h\iint\max\left(\frac{\partial}{\partial t}z(t)\psi(x,y)\right)^{2}\,\mathrm{d}x\mathrm{d}y$
@@
We can similarly obtain the effective spring constant, $k_{\mathrm{eff}}=\Omega_{0}^{2}m_{\mathrm{eff}}$, from the elastic potential energy:
@@
In this derivation, Eq. (25) uses the two-dimensional version of the potential energy density in a string [24]. Equation (26) can be interpreted as the fact that the tensile restoring force acts to reduce deformation (strain) in the membrane. A mode with greater total strain therefore has a stronger effective spring constant.
@@
This equation constrains $\psi$ and provides some intuition for the relationship between modeshape and eigenfrequency. The dimensionless numerator scales with increasing strain and the denominator increases with mode area. This corresponds to the fact that eigenmodes confined to smaller areas with greater strain have higher eigenfrequencies.
@@ IV.3 Damping
$m_{\mathrm{eff}}\frac{\mathrm{d}^{2}z(t)}{\mathrm{d}t^{2}}+\Gamma_{\mathrm{eff}}\frac{\mathrm{d}z(t)}{\mathrm{d}t}+k_{\mathrm{eff}}z(t)=F(t).$ (28)
Here, $\Gamma_{\mathrm{eff}}$ is the effective damping force coefficient, $F(t)$ is the drive force. It is worth reiterating that all of the effective quantities are defined with respect to a single eigenmode.
@@ IV.4 Spectral response and quality factor
$m_{\mathrm{eff}}\frac{\mathrm{d}^{2}z(t)}{\mathrm{d}t^{2}}+m_{\mathrm{eff}}\gamma\frac{\mathrm{d}z(t)}{\mathrm{d}t}+k_{\mathrm{eff}}z(t)=F_{0}e^{i\Omega_{d}t}.$ (33)
@@
$\arg(Z_{0})=\arctan\left(\frac{\frac{\gamma}{\Omega_{0}}\frac{\Omega_{d}}{\Omega_{0}}}{\left(\frac{\Omega_{d}}{\Omega_{0}}\right)^{2}-1}\right).$ (35)
@@
The quality factor also equals the ratio of total energy stored in the oscillator, $E$, to the energy lost per oscillation, $\Delta E$:
@@
Eqs. (37) and (38) are equivalent in the small damping limit, where we can use Eq. (32) to calculate the ratio of total energy to energy lost per cycle:
$\frac{E}{\Delta E}$ $=\frac{z(0)^{2}e^{-\gamma t}}{z(0)^{2}(e^{-\gamma t}-e^{-\gamma(t+2\pi/\Omega)})}$
$=\frac{1}{1-e^{-2\pi\gamma/\Omega}}$
$\simeq\frac{\Omega}{2\pi\gamma},$ (39)
where in the last step we used the Taylor expansion under the justification that $\gamma/\Omega\ll 1$. One need only then substitute the ratio of energies into Eq. (38) to obtain Eq. (37).
@@
Figure 10: Experimental demonstration of how resonances correspond to distinct eigenmodes, using the example of a phononic error correcting device [69]. (a) Optical microscope image of a phononic error-correcting device which consists of three coupled silicon nitride resonators. The resonators are $100\,\mathrm{\upmu m}$ diameter circular drums, coupled with evanescent tunnels. The device in this image has tunnels of length $20\,\mathrm{\upmu m}$. (b) Spectral response from optical Doppler interferometry, measured on the bottom left resonator when driving on the top resonator. These data were taken on a device with $11\,\mathrm{\upmu m}$ tunnels. (c) Experimentally obtained images of the modeshapes labelled in subfigure (b), on the same device and again driving the top resonator. These images were taken by measuring the acoustic power spectral density around the drive frequency while driving at a particular resonance frequency. Higher-order drum modes can be identified by the number and position of their nodal lines (locations where the amplitude of vibration is zero). For instance, mode A corresponds to the $(m,n)=(0,1)$ circular drum mode, while mode D corresponds to the $(m,n)=(1,2)$ mode.
In this section we have considered just one resonance with effective parameters $k_{\mathrm{eff}}$, $m_{\mathrm{eff}}$, and $\gamma$. In practice a membrane will usually exhibit a range of resonances corresponding to the different eigenmodes. Fig. 10 shows an experimental example of this, where on a single device we can measure a large number of resonances, each corresponding to a particular spatial mode that can be experimentally driven and observed [69]. In Fig. 10(b) the resonance peak heights and widths vary because the different modes are more or less efficiently actuated by the electrode, and have larger or smaller damping rates. Differences in actuation efficiency can be observed for the mode labelled ‘C’ in Fig. 10(c). Section VII goes into more detail about how the mode shape affects actuation efficiency and loss.
@@ V Duffing nonlinearity
Nonlinearity enables many of the more useful and complex behaviours in phononic devices—for example it is essential for solitons and frequency combs [70, 71, 72, 73, 74], logic [19, 18, 75, 76, 20], and other phenomena such as four-wave mixing [77, 36, 78, 18] and bifurcations [79, 70]. In suspended membranes two key sources of nonlinearity are elongation stresses during large out-of-plane vibration, and electrostatic actuation nonlinearities. In both cases we can effectively model the nonlinearity by adding a quartic potential to the mechanical potential energy (we show this later). This model is called a Duffing nonlinearity, named after Georg Duffing (1861-1944) [80, 81].
@@
where $\alpha_{\mathrm{eff}}$ is the Duffing coefficient. Fig. 11 plots this model potential. The case $\alpha_{\mathrm{eff}}>0$ corresponds to a stiffening nonlinearity, which increases the wall steepness. The case $\alpha_{\mathrm{eff}}<0$ corresponds to a softening nonlinearity, which decreases the wall steepness.
@@
With Duffing nonlinearity the resonance frequency depends on the amplitude of oscillation. For a stiffening nonlinearity ($\alpha_{\mathrm{eff}}>0$) the membrane will increase in resonance frequency at larger amplitudes. For a softening nonlinearity ($\alpha_{\mathrm{eff}}<0$) the membrane will decrease in resonance frequency at larger amplitudes.
The shifting resonance frequency of a Duffing oscillator has widespread implications. It limits the dynamic range of a frequency-shift based nanomechanical sensor [82], and it complicates the problem of coupling multiple nonlinear oscillators [83]. It also introduces new functionalities, such as amplitude bistability which we will look at shortly in Section V.2.
To incorporate Duffing nonlinearity into the lumped element model, we note from Eq. (40) that the restoring force $F_{\mathrm{res}}$ is altered from Hooke’s law to $F_{\mathrm{res}}=-\mathrm{d}U/\mathrm{d}z=-k_{\mathrm{eff}}z-\alpha_{\mathrm{eff}}z^{3}$. Substituting the additional term into the lumped element dynamical equation (Eq. (28)) with a sinusoidal drive, we obtain the Duffing equation:
@@
Equation (45) quantifies the response of the resonator to the applied driving force. Figure 12(a) illustrates this response. As predicted in Eq. (42), we see the resonance frequency shifts depending on the amplitude of vibration, with the direction of the shift corresponding to the sign of $\alpha_{\mathrm{eff}}$. The peak response increases for softening nonlinearities ($\alpha_{\mathrm{eff}}<0$), because to first order a softening nonlinearity reduces the effective spring constant of the oscillator (even as the force $F_{0}$ remains constant).